OpenStudy (anonymous):

find the slope and 7-intercept of 3x-2y=24 -3x 2y=-3x+24 y=-3x+24/2 y=-3x+12 (0,-12) is the slope is better put as -3/-12 or simplified as 1/4

OpenStudy (anonymous):

thanks in advance

OpenStudy (aroub):

wait..so this is the question 3x-2y=24 , (0,-12) and its parallel??

OpenStudy (anonymous):

oh snap, no please disregard (01,-12) i made a mistake right there trying to interpret the y intercept

OpenStudy (anonymous):

(0,-12) argh

OpenStudy (anonymous):

slope is -3, y int is -12

OpenStudy (aroub):

-2y==3x+24 y=-2/3x-12 slope=-2/3 y-intercept=-12